# Larimar: Large Language Models with Episodic Memory Control

### v1

@@ 5.4 Selective Forgetting
Figure 2: Batch editing accuracy on counterfact dataset. Baseline performances are taken from [Meng et al., 2023]. Green: MEMIT, Orange: ROME, Magenta: MEND, Black: Larimar-6B.
The goal of this section is to check if a specific fact can be selectively erased from $N$ facts that are have been written in Larimar’s memory. We first checked if many edits can be written at once to memory and accurately retrieve from it. Figure 2 shows that the rewrite accuracy is near 100% for up to 512 edits (the memory size $K$) and then drops to 82% for 1024 edits. This result shows Larimar’s ability to compress more than $K$ facts into its size-$K$ memory. This performance level is higher when compared to baselines like MEND and ROME, but subpar compared to MEMIT [Meng et al., 2023], which can accurately handle a very large batch of edits at a cost of reduced editing speed (see Table 7) and is also not meant to handle sequential editing. Note that Larimar’s recall matches MEMIT for $N<K$ facts, and $K$ can be chosen as needed during inference.
To test the ability of Larimar for selectively forgetting specified facts during inference, we first write $N$ facts to memory ($\alpha_{i}=1$ in Eq. (4)), and then forget one fact ($\alpha_{i}=-1$), and also write to memory in its place (with $\alpha_{i}=1)$ the same fact with the answer replaced with the string “unknown.” We compare recall for the forgotten fact before and after the forgetting operation. To demonstrate that forgetting does not compromise other memories, we also report the recall on the remaining $N-1$ facts in memory. The samples used are from the ZsRE validation set and from the Counterfact test set. Table 4 reports these results, comparing to a $k$-shot in-context learning (see Appendix for example prompt) baseline with Llama2-13B, and showing that Larimar can selectively forget using the memory updating mechanism, while retaining the remaining knowledge, whereas in-context learning struggles.
Figure 3: Mean F1 score on a held-out set of unseen rephrasings from ZsRE over a sequence of 3000 edits, showing Larimar’s better generalization over GRACE on two datasets with $1000$ or $511$ independent facts ($10$ and $\approx 20$ rephrasings per fact, respectively).
@@
Attack Success (%) Larimar (single) 17.6 Larimar (batch) 21.5 ROME (single) 29.0 MEMIT (single) 49.3
Table 5: Input rephrasing attack success for a budget of 20 on Counterfact samples. Writing [prompt + ”unknown”] to Larimar-6B’s memory in single fact or batch mode is more effective in preventing answer leakage in generations, when compared to updating the decoder with an empty model response objective for a single fact via direct model editing methods on GPT-J 6B.
We also evaluate Larimar to prevent generation of specific information by writing an empty (i.e. censored) response for the corresponding prompt to the memory. The baselines we consider are ROME and MEMIT, which were adapted to delete information in [Patil et al., 2023]. Specifically, the authors updated the decoder $d$ with an empty response objective for a given string $x$ that is known to the decoder and is aimed to be deleted, such that the probability of an “empty” target string $E$ is maximized, $\text{argmax}_{d}\mathbb{P}[E|x,d]$. To probe into the effectiveness of information deletion, they used a blackbox input rephrasing attack, where the presence of information of interest was checked in a number of model outputs as the model was prompted with different input rephrases. For Larimar, a single input prompt followed by “unknown” (= empty response) is written to the memory during inference to prevent leakage of the answer in the decoded outputs. The attack is considered successful on a given input prompt if the answer is found within a fixed number of model generations obtained using prompt rephrases. Samples from Counterfact known to GPT-J 6B and Larimar’s decoder were used for this experiment. We used 5 sample responses for each of 4 paraphrases per fact (total attack budget 20), which were generated as prescribed in [Patil et al., 2023]. Table 5 shows the results, suggesting that writing to Larimar’s memory is more effective than direct model editing methods for preventing answer leakage for a single input prompt (17.6% attack success for Larimar, vs. 29% and 49% for ROME and MEMIT, respectively). Larimar can further be used to restrict the response for a batch of facts in one shot – the robustness to rephrase attacks remains still higher than baselines.
@@ 5.5 Generalization to long input context
We perform fact recall with long context using data that is not present in the base decoders pretraining corpus. For this purpose, we curated facts from CNN Fast Facts [CNN, 2023] for 2021, 2022 and 2023.
We divide the input text into $T$ chunks, which is in the range of Larimar’s training context window, and store each of these chunks in a separate memory $M_{i},i=1..T$. Given a query, we address and read from each of these memories. The readouts from these memories then form the basis of the successive memory, which is then queried and read from again. This process is continued until the number of readout in the final memory is similar to Larimar’s input training context window. This recursive search in the latent memory space and using readouts to construct new higher-level memory is performed to process the long context with Larimar’s memory trained on a relative small episode length. The retrieved $Z_{r}$ from the final successor memory is passed to the decoder to predict the response. It should be noted that memory hierarchy is also found in hippocampus and thought to be implicated in learning [Collin et al., 2015].
Table 6 shows Larimar’s recall performance does not degrade much with increasing input context length, even compared to some of most competitive baseline LLMs trained with longer training context. We also compare with Supersizing Transformer [Klett and Ahle, 2023] which is a memory-augmented model, however it did not show competitive recall performance because it was not trained to perform memory-conditioned generation. Due to memory processing in the latent space, Larimar is also efficient is terms of number of KV cache token computation compared to baseline methods. Our experiments on 128 facts case shows that the average time required by Larimar to read from memory is 0.36s compared to 1.44s for Mistral7b base model.
@@
Method Train Context $n_{fact}=64$ $n_{fact}=96$ $n_{fact}=128$ $n_{fact}=256$ mistral-7b (3-shot) 8192 0.98 / 2655 0.96 / 3495 0.57 / 4334 0.42 / 7417 gpt-neox-20b (3-shot) 2048 0.52 / 2366 0.36 / 3193 0.33 / 4020 0.35 / 7231 llama2-13b (3-shot) 4096 0.97 / 2755 0.66 / 3628 OOM OOM Supersizing Transformer 2048 0.39 / 1462 0.39 / 2249 0.37 / 3072 0.37 / 6201 Supersizing Transformer + filtering 2048 0.72 / 1640 0.71 / 2375 0.70 / 3110 0.69 / 5809 Larimar-1.3b 384/2048 0.89 / 1565 0.88 / 2276 0.88 / 2988 0.86 / 5607 Larimar-6b 384/2048 0.82 / 1565 0.81 / 2276 0.81 / 2988 0.80 / 5607
Table 6: Novel fact addition recall rate on FastFacts. Larimar shows good recall performance and can extrapolate to higher context length trained than it was trained on. Baseline models show good recall on small context but recall degrades significantly for higher context.
@@ 6 Related work
Transformers struggle with accessing and updating long-term memory [Fan et al., 2021]. Efforts to extend input context length for better performance encounter issues integrating inherent model knowledge with external facts, lacking robustness [Li et al., 2022, Liu et al., 2023]. Augmenting transformers with external, non-differentiable memory and k-nearest neighbor (kNN) attention has shown promise in improving language modeling by utilizing additional context [Grave et al., 2017, Khandelwal et al., 2019]. However, kNN-augmented models face challenges in controlling memory during decoding, leading to difficulties in updating facts due to conflicts between encoded knowledge and real-time information [Li et al., 2022, Liu et al., 2023, Zhu et al., 2020].
@@
For comprehensive surveys of editing approaches see [Yao et al., 2023, Zhang et al., 2024, Wang et al., 2023b]. Editing methods can be broadly categorized into three categories: ‘Recognition Phase’, ‘Association Phase’ and ‘Mastery Phase’  [Zhang et al., 2024]. The ‘recognition phase’-targeting methods consider demonstrating right context to help the LLM output correct facts, either via in-context demonstrations of similar examples [Zheng et al., 2023], or training an external model on edits [Mitchell et al., 2022].The ‘association phase’ -related editing methods consider merging new knowledge to that of the base LLM, either by patching (adding and training) error-specific neurons [Huang et al., 2023], or by adding a an adaptor storing edit key-value pairs to a specific LLM layer [Hartvigsen et al., 2022]. The ‘mastery phase’ methods learn to update base LLM’s own parameters. Examples are regularized finetuning [Zhu et al., 2020] and hypernetwork-based methods [Mitchell et al., 2021, De Cao et al., 2021]. Recent works also explore the ‘locate-then-edit’ approach: [Meng et al., 2022a, b] first perform a causal tracing to detect which part of hidden states can be attributable to the fact and then do a rank-one update of the corresponding weight parameters to directly write in the updated fact.
@@
Current model editing approaches, while promising [Yao et al., 2023], face significant limitations, such as high training costs and difficulties in generalizing to new data . These methods often cannot efficiently update Large Language Models (LLMs) due to extensive time and memory requirements [Mitchell et al., 2022]. Furthermore, the assumption that knowledge within LLMs is localized has been challenged [Hase et al., 2023], indicating that simple parameter updates may not be effective for comprehensive edits. The performance of LLMs degrades with multiple edits, leading to issues like knowledge forgetting and distortion [Mitchell et al., 2022, Meng et al., 2023, Gupta et al., 2024, Li et al., 2023, Gu et al., 2024]. Alternatives like external cache or memory-based editing have been proposed to circumvent direct model modifications, yet challenges in selectively forgetting outdated or sensitive knowledge persist [Ishibashi and Shimodaira, 2023, Patil et al., 2023].
@@
Different from the above-mentioned works, we present a novel approach to augment Large Language Models (LLMs) with generative memory, enabling dynamic editing and adaptation without retraining. This differs from traditional methods that update LLM parameters [Meng et al., 2022a, b] or external memories [Han et al., 2023, Hartvigsen et al., 2022] and does not require multiple demonstrations for control Zheng et al. [2023].
@@
Larimar’s forgetting operation does not use negative examples to fine-tune LLMs for unlearning [Yu et al., 2023]. Neither Larimar requires tailored fine-tuning [Eldan and Russinovich, 2023] or inserting extra layers [Chen and Yang, 2023], and is complimentary to in-context unlearning approaches such as [Pawelczyk et al., 2023] for fact forgetting.
@@ 7 Conclusions
In this work, we propose augmenting LLMs with a dynamically updatable and distributed episodic memory as a means to online knowledge adaptation. By exploiting a one-shot memory update mechanism, combined with memory-conditioned decoding, the proposed framework shows accurate, precise, robust, and significantly faster editing performance compared to baselines in single-fact, as well as the challenging sequential editing experiments. We exploit the same memory updating mechanism to enable a fast and selective fact forgetting operation, as well as an effective information deletion mechanism. We also provide a simple approach for handling long input context by recursively reading from Latimar’s memory space, revealing better fact recall from long input context by Larimar when compared to state-of-the-art LLMs trained with a much larger training context window. The proposed framework thus provides a simple, general, and principled approach to update LLMs in real-time by coupling them with an adaptable episodic memory control.
8 Broader Impact and Ethical Considerations
@@ Appendix A Baselines
MEMIT MEMIT aims direct model editing via fact tracing and followed by parameter editing. It is an expanded version of ROME, which enables the editing of large amounts of factual data through the updating of a sequence of MLP layers.
@@ A.1 Prompt Example For Forgetting with ICL experiments
Appendix B Ablation Results on CounterFact dataset
In Table 8 we show that when Larimar has access to additional fact paraphrases, its paraphrase performance increases from 76.5 to 82.8. Note that in this setup the average number of added paraphrased facts is one and we queried the model with paraphrased prompts unseen by the memory. Also, observe that the use of the scope detector for query detection is crucial for the model’s performance to properly handle the neighborhood prompts.
Editor Edit Success Paraphrase Neighborhood Larimar-6B w/ scope 99.6 76.5 80.2 Larimar-6B +para 99.6 82.8 80.6 Larimar-6B +para, no scope 99.6 88.7 16.3
Table 8: Single fact edit valuation on CounterFact dataset. Larimar-6B base is the baseline which includes only a single fact in the memory and uses in-scope query detector. Larimar-6B +para is the version which adds into the memory on average one additional paraphrased fact.
@@ Appendix B Ablation Results on CounterFact dataset
Config Editor Metrics Edit Success Paraphrase Neighb S M S M S M Larimar 100.0 99.7 38.4 -2.9 74.2 1.6 C1 No Scope 100.0 99.8 37.8 -3.0 22.4 -34.1 No Memory 23.3 -4.4 26.5 -3.5 77.7 4.7 Larimar 100.0 99.9 35.2 -3.5 75.4 2.0 C2 No Scope 100.0 99.9 33.1 -3.6 26.2 -36.2 No Memory 20.6 -4.9 24.5 -4.1 78.9 5.4 Larimar 100.0 99.8 41.9 0.4 74.8 1.6 C3 No Scope 100.0 99.9 41.1 0.4 14.3 -45.8 No Memory 21.6 -4.8 25.4 -3.8 78.4 5.0
@@
Config Editor Metrics Edit Success Paraphrase Neighb S M S M S M Larimar 99.6 96.0 76.3 22.1 80.2 3.9 C1 No Scope 99.6 96.1 83.6 23.6 10.4 -32.8 No Memory 15.8 -6.8 18.6 -6.8 83.6 6.9 Larimar 42.4 3.4 37.8 -2.7 82.9 6.9 C2 No Scope 42.5 3.4 38.8 -2.7 67.8 5.7 No Memory 15.2 -7.0 18.3 -6.3 83.2 6.9 Larimar 99.9 98.9 68.0 10.8 79.9 3.1 C3 No Scope 99.9 99.0 81.3 11.9 7.2 -48.1 No Memory 15.0 -6.6 18.5 -6.2 83.6 6.5 Larimar 91.1 70.8 66.1 16.1 81.6 5.9 C4 No Scope 91.0 70.8 67.9 16.3 27.8 -5.4 No Memory 15.4 -6.9 18.1 -6.0 83.2 6.6 Larimar 99.9 98.9 72.2 15.1 79.7 3.3 C5 No Scope 99.9 99.0 82.9 16.5 6.6 -50.4 No Memory 14.3 -6.9 18.8 -6.3 83.5 6.8
@@
Editor Edit Success Paraphrase Neighborhood S M S M S M Larimar (ESD) 99.6 96.0 76.3 22.1 80.2 3.9 Larimar(ISD-ep4) 99.8 89.8 83.5 23.5 82.1 6.1 Larimar(ISD-ep8) 99.9 92.0 82.9 16.5 81.3 5.3
@@
Appendix C ZsRE Single Fact Editing Results

### v2

@@ 5.4 Selective Forgetting
Figure 2: (a) Batch editing accuracy on Counterfact dataset. Baseline performances are taken from [Meng et al., 2023]. Green: MEMIT, Orange: ROME, Magenta: MEND, Black: Larimar-6B. (b) Mean F1 score on a held-out set of unseen rephrasings from ZsRE over a sequence of 3000 edits, showing Larimar’s generalizes better over GRACE on two datasets with $1000$ or $511$ independent facts ($10$ and $\approx 20$ rephrasings per fact, respectively).
The goal of this section is to check if a specific fact can be selectively erased from a batch of $N$ facts that are written to Larimar’s memory in one-shot. We first checked the batch editing performance of Larimar. Figure 2 (a) shows that the rewrite accuracy is near 100% for up to 512 edits (eqv. to the memory size $K$) and then drops to 82% for 1024 edits. This result shows Larimar’s ability to effectively compress more than $K$ facts into its size-$K$ memory (Figure 3 in appendix). This performance level is higher when compared to baselines like MEND and ROME, but subpar compared to MEMIT [Meng et al., 2023], which can accurately handle a very large batch of edits at a cost of reduced editing speed (see Table 7) and is also not meant to handle sequential editing. Note that Larimar’s recall matches MEMIT for $N<K$ facts, and $K$ can be chosen as needed during inference (Figure 4 in appendix).
To test the ability of Larimar for selectively forgetting specified facts during inference, we first write $N$ facts to memory ($\alpha_{i}=1$ in Eq. (4)), and then forget one fact ($\alpha_{i}=-1$), and also write to memory in its place (with $\alpha_{i}=1)$ the same fact with the answer replaced with the string “unknown.” We compare recall for the forgotten fact before and after the forgetting operation. To demonstrate that forgetting does not compromise other memories, we also report the recall on the remaining $N-1$ facts in memory. The samples used are from the ZsRE validation set and from the Counterfact test set. Table 4 reports these results, comparing to a $k$-shot in-context learning (see Appendix) baseline with Llama2-13B, and showing that Larimar can selectively forget using the memory updating mechanism, while retaining the remaining knowledge, whereas in-context learning struggles.
@@
ROME (s) MEMIT (s) Larimar (s) Larimar (b) Attack Success (%) 29.0 49.3 17.6 21.5
Table 5: Input rephrasing attack success: Larimar-6B in-memory writing (single fact (s) or batch (b) mode) vs. GPT-J 6B editing.
We also evaluate Larimar to prevent generation of specific information by writing an empty (i.e., censored) response for the corresponding prompt to the memory. The baselines we consider are ROME and MEMIT, which were adapted to delete information in [Patil et al., 2023]. Specifically, the decoder $d$ was updated with an empty response objective for a given string $x$ that is known to the decoder and is aimed to be deleted, such that the probability of an “empty” target string $E$ is maximized, $\text{argmax}_{d}\mathbb{P}[E|x,d]$. A blackbox input rephrasing attack was then used; the presence of information of interest was checked in a number of model outputs as the model was prompted with different input rephrases. For Larimar, a single input prompt followed by “unknown” (= empty response) is written to the memory during inference to prevent leakage of the answer in the decoded outputs. The attack is considered successful on a input prompt if the answer is found within a fixed number of model generations obtained using prompt rephrases. About 300 samples from Counterfact known to GPT-J 6B and Larimar’s decoder were used for this experiment. We used 5 sample responses for each of 4 paraphrases per fact (total attack budget of 20), which were generated as prescribed in [Patil et al., 2023]. Table 5 shows the results, suggesting that writing to Larimar’s memory is more effective than direct model editing methods for preventing answer leakage for a single input prompt (17.6% attack success for Larimar, vs. 29% and 49% for ROME and MEMIT, respectively). Larimar can further restrict the response for a batch of facts in one shot – the robustness to rephrase attacks remains still higher than baselines.
@@ 5.5 Generalization to long input context
We perform fact recall with long context using data that is not present in the base decoders pretraining corpus. For this purpose, we curated facts from CNN Fast Facts [CNN, 2023] for 2021, 2022, and 2023. We divide the input text into $T$ chunks, which is in the range of Larimar’s training context window, and store each of these chunks in a separate memory $M_{i},i=1..T$. Given a query, we address and read from each of these memories. The readouts from these memories form the basis of the successive memory, which is then queried and read from again. This process is continued until the number of readout in the final memory is similar to Larimar’s input training context window. The recursive search in latent memory space and using readouts to construct new higher-level memory is performed to process the long context with Larimar’s memory trained on a relative small episode length. The retrieved $Z_{r}$ from the final successor memory is passed to the decoder for predicting response. It should be noted that memory hierarchy is found in hippocampus and is implicated in learning [Collin et al., 2015].
Table 6 shows Larimar’s recall does not degrade much with increasing input context length, even compared to some of most competitive baseline LLMs trained with longer training context. We also compare with Supersizing Transformer [Klett and Ahle, 2023], a memory-augmented model, however it did not show competitive recall performance because it was not trained to perform memory-conditioned generation. Due to memory processing in the latent space, Larimar is also efficient is terms of number of KV cache token computation compared to baseline methods. Our experiments on 128 facts show that the average time required by Larimar to read from memory is 0.36s compared to 1.44s for Mistral-7b base model.
@@
Method Train Context $n_{fact}=64$ $n_{fact}=96$ $n_{fact}=128$ $n_{fact}=256$ mistral-7b (3-shot) 8192 0.98 / 2655 0.96 / 3495 0.57 / 4334 0.42 / 7417 gpt-neox-20b (3-shot) 2048 0.52 / 2366 0.36 / 3193 0.33 / 4020 0.35 / 7231 llama2-13b (3-shot) 4096 0.97 / 2755 0.66 / 3628 OOM OOM Supersizing Transformer 2048 0.39 / 1462 0.39 / 2249 0.37 / 3072 0.37 / 6201 Supersizing Transformer + filtering 2048 0.72 / 1640 0.71 / 2375 0.70 / 3110 0.69 / 5809 Larimar-1.3b 384/1024 0.89 / 1565 0.88 / 2276 0.88 / 2988 0.86 / 5607 Larimar-6b 384/2048 0.82 / 1565 0.81 / 2276 0.81 / 2988 0.80 / 5607
Table 6: Novel fact addition recall rate on FastFacts. Larimar shows good recall performance and can extrapolate to higher context length than it was trained on. Baseline models show good recall on shorter context but recall degrades significantly for longer context.
@@ 6 Related work
Transformers struggle with accessing and updating long-term memory [Fan et al., 2021]. Efforts to extend input context length struggle integrating inherent model knowledge with external facts, thereby lacking robustness [Li et al., 2022, Liu et al., 2023]. Augmenting transformers with external, non-differentiable memory and k-nearest neighbor (kNN) attention has shown promise in improving language modeling by utilizing additional context [Grave et al., 2017, Khandelwal et al., 2019]. However, kNN-augmented models face challenges in controlling memory during decoding, leading to difficulties in updating facts due to conflicts between encoded knowledge and real-time information [Liu et al., 2023, Zhu et al., 2020].
@@
For comprehensive surveys of editing approaches see [Yao et al., 2023, Zhang et al., 2024, Wang et al., 2023a]. Editing methods can be broadly categorized into three categories: ‘Recognition Phase’, ‘Association Phase’ and ‘Mastery Phase’  [Zhang et al., 2024]. The ‘recognition phase’-targeting methods consider demonstrating right context to help the LLM output correct facts, either via in-context demonstrations of similar examples [Zheng et al., 2023], or training an external model on edits [Mitchell et al., 2022].The ‘association phase’ -related editing methods consider merging new knowledge to that of the base LLM, either by patching (adding and training) error-specific neurons [Huang et al., 2023], or by adding a an adaptor storing edit key-value pairs to a specific LLM layer [Hartvigsen et al., 2022]. The ‘mastery phase’ methods learn to update base LLM’s own parameters. Examples are regularized finetuning [Zhu et al., 2020] and hypernetwork-based methods [Mitchell et al., 2021, De Cao et al., 2021]. Recent works also explore the ‘locate-then-edit’ approach: [Meng et al., 2022a, b] first perform a causal tracing to detect which part of hidden states can be attributable to the fact and then do a rank-one update of the corresponding weight parameters to directly write in the updated fact.
@@
Current model editing approaches, while promising [Yao et al., 2023], face significant limitations, such as high training costs and difficulties in generalizing to new data. These methods often cannot efficiently update LLMs due to extensive time and memory requirements [Mitchell et al., 2022]. Furthermore, the assumption that knowledge within LLMs is localized has been challenged [Hase et al., 2023], indicating that simple parameter updates may not be effective for comprehensive edits. The performance of LLMs degrades with multiple edits, leading to issues like knowledge forgetting and distortion [Mitchell et al., 2022, Meng et al., 2023, Gupta et al., 2024, Li et al., 2023, Gu et al., 2024]. Alternatives like external cache or memory-based editing have been proposed to circumvent direct model modifications, yet challenges in selectively forgetting outdated or sensitive knowledge persist [Ishibashi and Shimodaira, 2023, Patil et al., 2023].
@@
Different from the above-mentioned works, we present a novel approach to augment LLMs with generative memory, enabling dynamic editing and adaptation without retraining. This differs from existing works that update LLM parameters [Meng et al., 2022a, b] or external memories [Han et al., 2023, Hartvigsen et al., 2022], or requires multiple in-context demonstrations [Zheng et al., 2023].
@@
Larimar’s forgetting operation does not use negative examples to fine-tune LLMs for unlearning [Yu et al., 2023]. Neither Larimar requires tailored fine-tuning [Eldan and Russinovich, 2023] or inserting extra layers [Chen and Yang, 2023], and is complimentary to in-context unlearning (e.g., [Pawelczyk et al., 2023]) for fact forgetting.
@@ 7 Conclusions
In this work, we propose augmenting LLMs with a dynamically updatable and distributed episodic memory as a means to online knowledge adaptation. By exploiting a one-shot memory update mechanism, combined with memory-conditioned decoding, the proposed framework shows accurate, precise, robust, and significantly faster editing performance compared to baselines in single-fact, as well as the challenging sequential editing experiments. We exploit the same memory updating mechanism to enable a fast and selective fact forgetting operation, as well as an effective information deletion mechanism. We also provide a simple approach for handling long input context by recursively reading from Larimar’s memory space, revealing better fact recall from long input context by Larimar when compared to state-of-the-art LLMs trained with a much larger training context window. The proposed framework thus provides a simple, general, and principled approach to update LLMs in real-time by coupling them with an adaptable episodic memory control.
One obvious limitation of current Larimar architecture is being able to only handle shorter-lengh facts. In addition, the training is currently limited to sentence completion tasks. Future works will include expanding Larimar to modeling longer sentences and to more tasks like question answering, and summarization. Further, we will subject Larimar to more challenging tasks during inference, whereas the knowledge update in latent memory is needed for the model to navigate the task, for example in conversational setting.
Impact Statement
@@ Appendix A Baselines
MEMIT
MEMIT aims direct model editing via fact tracing and followed by parameter editing. It is an expanded version of ROME, which enables the editing of large amounts of factual data through the updating of a sequence of MLP layers.
SERAC
SERAC is a retrieval-based editing algorithm which uses a retrieval-based component consisting of an external memory that contains an explicit cache of edits. In addition, an edit scope classifier and a counterfactual model are trained using these edits. If the new input is identified as within the scope, the output from the counterfactual model is returned. If not, the base model is used for generation.
@@
In Table 1, we report the wall clock time for a single edit (averaged over 10 edits) on the CounterFact dataset for ROME [Meng et al., 2022a] and GRACE [Hartvigsen et al., 2022] that were computed using the EasyEdit [Wang et al., 2023b] framework with a single A100 (80G) GPU. Below we provide wall clock time of different existing editing methods, as reported by [Yao et al., 2023].
@@ A.1 Prompt Example For Forgetting with ICL experiments
Appendix B Ablation experiments on CounterFact single fact editing
In Table 8 we show that when Larimar has access to additional fact paraphrases, its paraphrase performance increases from 88.4 to 93.6. Note that in this setup the average number of added paraphrased facts is one or two and we queried the model with paraphrased prompts unseen by the memory. Also, observe that the use of the scope detector for query detection is crucial for the model’s performance to properly handle the neighborhood prompts.
Editor Edit Success Paraphrase Neighborhood Larimar-6B w/ scope 99.6 88.4 80.4 Larimar-6B w/ scope + 1 rephrase 99.7 92.9 79.3 Larimar-6B w/ scope + 2 rephrases 99.8 93.6 79.2 Larimar-6B w/o Scope 99.6 93.6 13.7 Larimar-6B w/o Scope + 1 rephrase 99.7 95.9 11.0 Larimar-6B w/o Scope + 2 rephrases 99.8 96.2 10.5
Table 8: Single fact edit valuation on CounterFact dataset. Larimar-6B w/ Scope is the baseline which includes only a single fact in the memory and uses in-scope query detector. Larimar-6B + rephrase is the version which adds into the memory on average one or more additional paraphrased facts during test and queries the memory with an unseen rephrased prompt. Results without a scope detector (w/o Scope) are also reported.
@@ Appendix B Ablation experiments on CounterFact single fact editing
Config Editor Metrics Edit Success Paraphrase Neighb S M S M S M Larimar 100.0 99.7 81.3 48.9 75.5 2.1 C1 No Scope 100.0 99.8 81.9 48.66 28.5 -27.4 No Memory 23.3 -4.4 26.5 -3.5 77.7 4.7 Larimar 100.0 99.9 80.3 49.1 74.7 1.9 C2 No Scope 100.0 99.9 80.3 51.2 24.5 -36.9 No Memory 20.6 -4.9 24.5 -4.1 78.9 5.4 Larimar 100.0 99.8 85.4 56.7 74.7 1.6 C3 No Scope 100.0 99.9 87.7 57.2 15.1 -46.1 No Memory 21.6 -4.8 25.4 -3.8 78.4 5.0
@@
In Table 11 we show ablation results with different scope detectors (detects whether a given prompt is related or not to the facts written in memory). ESD (externally trained scope detector) shows overall good performance across all three metrics (edit success, paraphrase and neighborhood). ISD (internal trained scope detector on train st of CounterFact dataset) shows some improvement in edit success and paraphrase while dropping the performance on neighborhood metric, which is expected as now it is less accurate in classifying the neighborhood prompts as out of scope.
@@
Config Editor Metrics Edit Success Paraphrase Neighb S M S M S M Larimar 99.6 96.0 87.6 53.9 80.6 4.4 C1 No Scope 99.6 96.1 94.6 55.5 15.7 -17.1 No Memory 15.8 -6.8 18.6 -6.8 83.6 6.9 Larimar 69.8 11.5 59.2 5.7 82.7 6.8 C2 No Scope 70.8 11.6 64.4 6.4 62.7 4.2 No Memory 15.2 -7.0 18.3 -6.3 83.2 6.9 Larimar 99.6 98.9 88.8 59.1 80.3 3.6 C3 No Scope 99.9 99.0 95.3 60.6 15.3 -22.1 No Memory 15.0 -6.6 18.5 -6.2 83.6 6.5 Larimar 91.0 69.7 78.9 34.7 81.6 5.9 C4 No Scope 91.0 69.7 83.9 35.4 29.8 -4.2 No Memory 15.4 -6.9 18.1 -6.0 83.2 6.6 Larimar 99.9 98.9 88.7 59.9 80.1 3.7 C5 No Scope 99.9 98.9 94.9 61.4 15.7 -22.8 No Memory 14.3 -6.9 18.8 -6.3 83.5 6.8
@@
Editor Edit Success Paraphrase Neighborhood S M S M S M Larimar (ESD) 99.6 95.9 87.6 25.9 80.6 4.3 Larimar(ISD-ep4) 99.6 96.6 94.6 55.5 15.7 -17.0 Larimar(ISD-ep8) 99.6 94.9 86.6 46.9 45.7 -5.9
@@
Appendix C ZsRE single fact editing experiments and ablations